OpenStudy (anonymous):

how do i solve 1/4 times √3 times 7√5^2???

OpenStudy (m):

you solve individually then multiply

OpenStudy (m):

\[\sqrt{3} = 3^{1/2}\]

OpenStudy (m):

you sort of need calculator for that

OpenStudy (m):

\[7\sqrt{5^{2}} = 7\sqrt{25} = 7(5)\]

OpenStudy (m):

= 35

OpenStudy (anonymous):

Gracias:)

OpenStudy (anonymous):

\[(1/4)*\sqrt{3}*7*\sqrt{5}{}^{\wedge}2=\frac{35 \sqrt{3}}{4} \]